Solve the problem involving the third law of Kepler's planetary motion with complete solution. It takes 365.25 days for Earth to complete one revolution around the sun. Calculate the A) period B) frequency C) angular frequency of Earth as it revolves around the sun


1
Expert's answer
2021-12-22T14:12:43-0500

A) We can find the period as follows:


"T=365.25\ days\times\dfrac{24\ h}{1\ day}\times\dfrac{3600\ s}{1\ h}=3.16\times10^7\ s."

B) We can find the frequency as follows:


"f=\dfrac{1}{T}=\dfrac{1}{3.16\times10^7\ s}=3.16\times10^{-8}\ Hz."

C) We can find the angular frequency as follows:


"\omega=2\pi f=2\pi\times3.16\times10^{-8}\ Hz=2.0\times10^{-7}\ \dfrac{rad}{s}."